### v1

# Automatic Intermodal Loading Unit Identification using Computer Vision: A Scoping Review

Abstract
The standardisation of Intermodal Loading Units (ILUs), such as containers, semi-trailers and swap bodies, has revolutionised global trade yet their efficient and robust identification remains a critical bottleneck in high-throughput ports and terminals. This paper reviews 63 empirical studies that propose computer vision (CV) based solutions. It covers the last 35 years (1990-2025), tracing the field’s evolution from early digital image processing (DIP) and traditional machine learning (ML) to the current dominance of deep learning (DL) techniques. While CV offers cost-effective alternatives for other types of identification techniques, its development is hindered by the lack of publicly available benchmarking datasets. This results in high variance for the reported results such as end-to-end accuracy ranging from 5 % to 96 %. Beyond dataset limitations, this review highlights the emerging challenges especially introduced by the shift from character-based text recognition to scene-text spotting and the integration of mobile cameras (e.g. drones, sensor equipped ground vehicles) for dynamic terminal monitoring. To advance the field, the paper calls for standardised terminology, open-access datasets, shared source code, while outlining future research directions such as contextless text recognition optimised for ISO6346 codes.
keywords:
intermodal loading unit identification , container code recognition , text detection , text recognition , logistics automation
journal: Transport Reviews
ILU
Intermodal Loading Unit
TU
transportation unit
UIC
International Union of Railways
UAV
unmanned aerial vehicle
IGD
InteGreatDrones
CER
character error rate
IoU
intersection over union
mAP
mean average precision
AP
average precision
FPS
frames per second
GFLOP
giga-floating point operations per second
ID
identifier
TRUDI
TRansportation Unit Detection and Identification
PRISMA-ScR
Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews
CMT
Containerised Maritime Transport
PCC
Population–Concept–Context
DIP
Digital Image Processing
ML
Machine Learning
DL
Deep Learning
AI
Artificial Intelligence
CV
Computer Vision
OCR
Optical Character Recognition
ACCR
Automatic Container Code Recognition
ALPR
Automatic Licence Plate Recognition
SVM
Support Vector Machine
MLP
Multi-Layer Perceptron
CNN
Convolutional Neural Network
GAN
Generative Adversarial Network
CCA
Connected Component Analysis
MSER
Maximally Stable Extremal Regions
SIFT
Scale-Invariant Feature Transform
HoG
Histogram of Gradients
LBP
Local Binary Pattern
[…]
[cv]organization=University of Hamburg,addressline=Department of Informatics, Computer Vision Group, city=Hamburg, country=Germany
[…]
[dos]organization=University of Hamburg,addressline=Department of Informatics, Distributed Operating Systems Group, city=Hamburg, country=Germany
[…]
[tuhh]organization=Hamburg University of Technology,addressline=Institute of Maritime Logistics, city=Hamburg, country=Germany
1 Introduction
The evolution of cargo transportation has been significantly transformed by the concept of containerisation. Before, cargo handling was labour-intensive, neither standardised nor suitable for automatisation. Handling and storage arrangements were costly and required special attention [1, 2], resulting in terminal congestion. Containerisation has facilitated automation and improved operational efficiency in cargo handling [3], while also significantly contributed to the growth of global trade [4]. The standardisation of freight handling has not only led to the shipping container but also led to a variety of standardised types of Intermodal Loading Units.
The term ILU has been introduced to describe units designed to be easily transferred between different modes of transport, such as vessels, trains, and trucks, without the need to unload and reload the contents. The loading units, such as containers, interchangeable bodies (swap bodies) and semi-trailers, fall under this term [5]. These units are typically built according to specifications defined in standards such as ISO668 [6]. These specifications ensure compatibility of ILUs with different transport systems [7] for smooth movement of goods along supply chains [8]. A unique identification code, following the ISO6346 [9] standard, is assigned to each ILU for tracking purposes. This code consists of four letters, a six-digit number and a check digit, which together indicate the owner and type of the ILU.
As global trade expands, many leading ports handle tens of millions of containers annually, with some operating at or above their nominal design capacity [10]. In this high throughput, keeping track of ILUs is important to ensure efficient management [11]. The often identically looking ILUs can exclusively be distinguished by their ID codes. Therefore, a reliable system that can rapidly recognise these codes is indispensable. Standardisation and digitalisation enabled the development of various techniques for ILU identification. Among these, Computer Vision (CV)-based identification code recognition stands out as a cost-effective alternative to RFID, which achieves only about 70 % accuracy and entails high installation and maintenance costs [12, 13]. In contrast, CV-based methods can be implemented using a range of camera setups from high-end static Optical Character Recognition (OCR) gates [14], that can identify entering and exiting ILUs [15, 16], to more affordable and flexible options such as using the existing CCTV infrastructure, cameras mounted on aerial or ground vehicles, or even mobile devices, which significantly reduce the need for specialised hardware. Furthermore, CV-based identification plays a crucial role in enhancing security by providing accurate verification for customs clearance, which can support in the correct declaration of goods [17, 18, 19].
Although CV-based methods reduce human-related errors and increase the processing speed [20, 21, 22, 23] there are several challenges: Containers often arrive in rapid succession, will be stacked for storage, and can be positioned at angles that partially occlude their ID codes. Environmental factors such as rain, fog, poor lighting, or glare from sunlight can degrade image quality, while harsh conditions, such as dirt, rust, or physical damage, further complicate automated ID code recognition [14, 24]. In addition, variations in font style, colour and placement of ID codes across different operators and regions create inconsistencies that reduce the reliability of automatic recognition systems.
At transfer points of the terminal to other modes of transport, each ILU must be identified and cross-checked with its associated documentation before entry or exit. When automated systems fail to achieve reliable recognition, manual verification becomes necessary, further slowing down the process. Even a brief slowdown at a busy gate can cause truck queues stretching hundreds of meters [25], disrupting yard operations and increasing turnaround times for transport vehicles. Extended waiting times contribute to congestion which increases fuel consumption and greenhouse gas emissions from idling trucks, amplifying the environmental footprint of port activities [26]. Therefore, efficient ILU identification is essential for overcoming these operational bottlenecks and sustaining the growth of containerised transport.
Despite its crucial role in improving transportation and trade security, to the best of our knowledge, there are not any literature reviews on ILU identification. To fill this gap, we prepared this literature review by following Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) guidelines [27] with the intention of identifying the available publications and proposed concepts.
In summary, there are 4 contributions of this review:
  • 1.
    Catalogue all empirical studies on Intermodal Loading Unit (ILU) identification using Computer Vision (CV) methods between 1990 and 2025.
  • 2.
    Re-evaluate the terminology used in ILU identification field and provide clear definitions.
  • 3.
    Synthesise an overview of Intermodal Loading Unit (ILU) identification methods, image acquisition settings, datasets, and evaluation metrics used in these studies.
  • 4.
    Highlight research gaps and suggest future priorities.
2 Terminology
This review distinguishes the roles of Digital Image Processing (DIP), Machine Learning (ML), and Deep Learning (DL) in the context of CV-based ILU identification to ensure precise use of technical terms and prevent misinterpretation across disciplines. Therefore, we provide the definitions of the general approaches:
Digital Image Processing (DIP): Classical, rule-based image operations (e.g., denoising [28], contrast normalisation [29], edge detection [30, 31], morphologic operations [32], thresholding [33, 34], geometric transforms [35]). These operations can be further extended for extracting hand-crafted features, such as edge detection [36, 30], Scale-Invariant Feature Transform (SIFT) [37] or Histogram of Gradients (HoG) [38]. It requires no training but is sensitive to illumination, motion blur, occlusions, or viewpoint changes [39, 40, 41]. The use cases of DIP methods in ILU identification are image enhancement [42], text localisation and character segmentation [43, 44], without relying on data-driven patterns.
Machine Learning (ML): While DIP methods rely on fixed, hand-crafted rules and deterministic operators (e.g. edge detectors, thresholding, morphologic filtering), ML introduces adaptability by learning from data. ML methods use features engineered via DIP (such as HoG, SIFT) as input to classifiers that can discriminate among classes (e.g. character classes, object vs. non-object, scene categories). Common models include Support Vector Machines [45, 46, 47], Random Forests [48], decision trees [49] and k-Nearest Neighbors [50]. These reduce manual rule-tuning and can generalize better under variations in viewpoint, scale, and lighting, but still depend critically on quality of the features and consistency in preprocessing [51].
Deep Learning (DL): A subfield of ML using deep neural networks, such as Convolutional Neural Networks [52] or transformers [53] that learn features and classifiers jointly and end-to-end. Although DL requires substantial training data and computational resources it automatise the feature engineering step. For similar tasks depending on similar features, deep learning methods can benefit from transfer learning [54] and fine-tuning [55].
Besides the terminology for general approaches, we would like to provide definitions for CV-specific terminology:
Object detection: Localising (drawing bounding boxes around) and classifying (assigning labels to) objects within an image. Unlike image classification (which labels are for the entire image), object detection handles multiple objects per image, providing both what and where the objects are [56, 57].
Instance segmentation: Localising (pixel-level regions, drawing polygons around) and classifying individual objects within an image. Unlike object detection, which uses bounding boxes, instance segmentation assigns pixel-level class labels and provides precise shapes of the objects [58].
Character segmentation: Localising and segmenting characters in an image, typically as a prior step for recognition [59].
Character recognition: Classifying a single already localised character into its alphanumeric label. This problem is often tackled as an image classification task where the input is a cropped character patch and the output is the label for this single character patch [60].
Character detection: Localising (drawing bounding boxes around) and classifying (assigning labels to) characters within an image. Similar to the object detection, character detection provides information about where individual characters are and which characters they are [61].
(Scene-)Text detection: Localising regions containing text in an image (e.g., street signs,) without extracting its textual content. Outputs bounding boxes (often quadrilaterals) around words or lines. Unlike character detection, it focuses on full words or lines, which makes it more robust against distractors [62].
(Scene-)Text recognition: Extracting textual information from detected text regions by converting them into machine-readable characters or words [63, 64].
Text spotting: End-to-end method that jointly detects and recognises text in a single step [65, 61].
2.1 Terminology Beyond OCR Gates
Initially, CV methods were applied at terminal gates, where the task is to automatically identify loading units as they entered or exited the facility. These gates are among the most vulnerable and critical points of a terminal, making reliable identification essential. The task performed there was to capture an image of the loading unit and read its identification code through DIP [44]. In the literature, this task is often referred to as Automatic Container Code Recognition (ACCR) [66, 67, 68, 69, 70, 71, 72] or transportation unit (TU) identification [73]. The focus of ACCR is on using CV to read the ISO6346 identification codes from containers at the point of entry. These entrances remained as the critical bottlenecks in terms of operational efficiency and congestion impact, so their optimisation becomes one of the key challenges in terminal management [74].
However, the term ACCR is not precise. First, the word container is too narrow, since not only containers but also semi-trailers and swap bodies carry ISO6346-compliant codes (as defined in EN 13044 [75]). Second, the word recognition does not fully capture the actual tasks involved. In computer vision, recognition usually means extracting text from a clean image, especially for the mobile cameras. The term recognition refers to extraction of text from an image in the CV field. However, this presupposes that the visual input image consists solely of the target text. This assumption is rarely valid in the real-world scenarios involving surveillance cameras or mobile cameras, where text is typically embedded within complex and cluttered scenes. In these cases, an additional stage, text detection is required to localise the text regions before recognition can be performed [76]. Over time, these methods were extended beyond the gates. With advancements in computer vision and robotics, identification is now performed at quay cranes [77], yard blocks [78], and movable objects such as reach stackers [79]. These new use cases introduce more variation and complexity than the original gate-based setting.
For these reasons, we propose the term Intermodal Loading Unit (ILU) identification as an inclusive designation for the task of determining which specific ILUs with ISO 6346–compliant ID codes are visible in an image. This term includes various loading units and also for the CV terminology such as ILU detection/segmentation stage introduced in recent papers [73, 80], scene-text detection and recognition. A term that covers all these stages aligned with CV terminology is important, as it would address shared challenges across the CV applications. For example, using a common terminology can link the relevant studies to the broader problems such as small object segmentation [81] or small text detection [82] which are actively studied in general purpose CV methods. Moreover, a common terminology can help in finding the distinct aspects of ILU identification. Unlike in most common text recognition needs, where a language model based encoder (e.g. TrOCR [83]) is useful due to contextual cues present in the natural language, the contextless nature of the target text in ILU identification makes this approach ineffective. Such models tend to overfit to frequently occurring text instances in training data [84], limiting their applicability in this context. Adopting the term ILU identification can provide conceptual clarity and establish a coherent basis for researchers and practitioners.
3 Methodology
We finalised our search for articles in August 2025, using Google Scholar and dblp-Computer Science Bibliography to find scientific articles on ILU identification using CV or DIP techniques. Our search terms included container code recognition, trailer code recognition, and identification of containers and trailers. As discussed in Section 2, the terminology used in this field is not standardised and it directly affects our search methodology. We checked the cited papers and ’related articles’ in Google Scholar and Research Rabbit to expand our pool of eligible articles.
As the first review on ILU identification, we applied inclusive eligibility criteria to capture the full scope of available evidence. These criteria were specified a priori before screening began. They follow the Population–Concept–Context (PCC) logic and the methodological guidance in [27]. Our review includes articles written in English, published from 1990 to 2025, that follow a scientific structure. Both peer-reviewed articles and preprints are included as long as they report quantitative results, allowing us to catalogue the latest research findings. The final selection of articles and data charting were determined through consensus among the authors, following a review of titles, abstracts, and full texts.
From the 63 eligible articles the following data items were extracted into a spreadsheet [^1] https://cloud.uni-hamburg.de/s/FHqaGLR4Ccy8ceG: journal or conference details, contributions, methods used, method class (DIP, ML, or DL), dataset specifics (availability, number of images, diversity in lighting and weather), image collection region and camera position such as fixed (CCTV) or moving (aerial or ground vehicle), evaluation metrics, results, deployment status, suggested future directions and funding. The resulting table was then analysed and discussed by the authors.
4 Results and Discussion
Our search yielded 91 articles published between 1990 and 2025. During the screening process, articles were excluded due to being duplicates ($n=2$), not reporting quantitative results ($n=6$), written in a non-English language ($n=5$), focusing solely on container detection or integrity assessment rather than identification via ISO6346 compliant codes ($n=6$), solutions for other industrial text recognition applications such as wagon code recognition ($n=5$), unavailability of their full text ($n=2$), or being patents ($n=2$).
Table 1: Eligible article characteristics ($n=63$). Five articles involve international collaborations, which results in the count of 69 rather than the 63, in geographic regions section.
Category Characteristic Count Percentage
Year of publication 1990–1995 2 3.17 %
1996–2000 1 1.59 %
2001–2005 5 7.94 %
2006–2010 3 4.76 %
2011–2015 10 15.87 %
2016–2020 18 28.57 %
2021–2025 24 38.10 %
Geographic region Asia 55 79.71 %
Europe 8 11.59 %
North America 3 4.35 %
Africa 2 2.90%
Oceania 1 1.45 %
Funding source Not reported 32 50.79 %
Public 24 38.10 %
Public-Industry Collab. 4 6.35 %
Industry 3 4.76 %
Rank Unranked 32 50.79 %
Q1 - Scimago 8 12.70 %
A - CORE 5 7.94 %
Q2 - Scimago 5 7.94 %
B - CORE 4 6.35 %
Q3 - Scimago 4 6.35 %
Q4 - Scimago 4 6.35 %
C - CORE 1 1.59 %
Table 1 presents the characteristics of the eligible articles. The data indicate a consistent increase in publications over time, especially from 2011 onwards. This trend reflects growing interest in automatic ILU identification systems, likely driven by advancements in CV as well as greater computational power affordable to major terminals and ports.
Asia contributes 79.71 % of the publications. This region handles high volumes of maritime traffic requiring efficient logistics and ILU identification systems. Countries such as China, Taiwan, South Korea, Vietnam, and Japan play significant roles in global trade, making effective logistics and supply chain management essential. Substantial investment in research and development by Asian governments and companies further supports this trend [85, 86].
Public funding is reported as the main source, supporting about two out of five of the publications, indicating strong governmental backing. This support is often motivated by the potential for technological advancements to benefit national infrastructure and economic growth [87, 88, 89]. Public funding is particularly important for the logistics sector that operates with low profit margins [90, 91]. Industry funding and public-industry collaborations constitute one ninth of the publications, suggesting that while there is some industry involvement, companies largely rely on public sector investments to drive technological development. However, the presence of pure industry funding indicates a genuine market demand for companies to invest in research to remain competitive [92, 11].
To analyse the rankings of journals and conferences we followed the rankings of Scimago [^2] https://www.scimagojr.com/ for journals and CORE [^3] https://www.core.edu.au/icore-portal for conferences. Table 1 shows that more than half of the articles on automatic ILU identification task are published in unranked outlets, which is not uncommon for applied or niche technical topics [93].
Among the ranked publications, Q1-Scimago and A-CORE have the highest numbers, suggesting that the topic is relevant for high-quality research venues. However, the majority of publications are spread across lower-tier or unranked venues. This distribution implies that the field could benefit from a greater emphasis on more novel contributions, increased comparability, and reproducibility of the results. This could be achieved through publicly available benchmark datasets and repositories that showcase the source code and the model weights. By adopting such research standards, ILU identification tasks with CV-based methods could be more relevant for high-ranked publication outlets.
Overall, the characteristics of the reviewed articles highlight a growing research field characterised by substantial activity in Asia and a predominance of public funding. Expanding funding sources and promoting industry collaboration may further advance automatic ILU identification through CV technologies.
4.1 Methods
Figure 1: Approach categories used for ILU identification task over the years.
The development of methods for ILU identification has undergone a significant transformation over three decades, as illustrated in Figure 1. For two decades between 1990 and 2010, the field relied predominantly on DIP techniques combined with early ML methods and to a lesser degree, DIP methods alone. This mirrors broader trends in the CV literature [94]. In this period, researchers mainly followed a three-stage approach: 1) Removal of the background using thresholding [95, 43], edge detection [43, 96] and morphological operations [97, 96]. 2) After segmenting the characters, each was classified by methods like template matching [98, 99, 100, 96], or traditional ML methods like Multi-Layer Perceptron (MLP) [101, 97, 44] or fuzzy logic [102]. 3) Finally, the predictions on the individual characters were postprocessed to output a full ID code. These methods performed adequately in controlled image acquisition systems like OCR-gates [103] but struggled with challenging conditions, such as variability of brightness [68] and damages [104, 105] on the ILU images.
In the subsequent five years, a modest increase in the use of ML can be observed, particularly in hybrid DIP and ML approaches [106, 107, 108, 109, 110, 12]. This reflects the growing recognition of ML’s potential for enhancing robustness in character and symbol recognition tasks [111] in CV. However, the field lagged several years behind the broad adoption of DL for other applications, despite its transformative influence on CV [112, 113] from 2012 onwards (e.g., the emergence of AlexNet [114]). This delay may be attributed to the scarcity of labelled ILU identification datasets, limited computational infrastructure, and the conservative approach for adopting new technologies in the logistics industry.
The five-year interval from 2016 to 2020 marks a turning point, with a notable adaptation of DL-based methods. Publications using only DL-based approaches rise sharply, as well as the hybrid strategies combining DIP, ML, and DL. This phase was driven by advances in GPU hardware [115, 116], and the availability of open-source frameworks such as TensorFlow [117] and PyTorch [118]. Deep architectures, particularly CNNs [119], rapidly proved their superiority for conducting ILU identification under challenging conditions [72, 120]. As a consequence, purely DIP-based approaches declined significantly, often limited to pre- or post-processing roles within hybrid pipelines [68, 80, 20, 121, 70].
ILU identification with fixed cameras often resembles OCR from scanned documents, as the background is mostly uniform. Segmenting and recognising characters is a sensible approach in such cases. Until 2016, all methods developed for ILU identification focused on character-level localisation and recognition, and this approach is still in use with DL methods [122, 123, 124]. However, with the increasing use of mobile cameras (e.g. vehicle-mounted cameras), the task has evolved to resemble scene-text detection and recognition. As a result of this, the trend is shifting towards using comprehensive scene-text detection and recognition methods rather than relying on character detection and recognition.
By 2021–2025, DL-based approaches had established themselves as the predominant paradigm for ILU identification. Even hybrid DIP-based methods virtually disappeared, indicating that the DL models now offer both the required accuracy and operational robustness [66, 125, 126]. The relative infrequency of hybrid approaches suggests that most effective solutions now rarely require DIP or traditional ML methods, except in specialised scenarios such as real-time constraints [104, 16, 18, 127, 124, 15].
4.2 Main Approaches for Deep Learning
Observing the growing interest of DL methods, we now focus on the CV tasks relevant for extracting textual information. OCR techniques are designed to identify and localize individual characters in an image. They are used predominantly for document text recognition and perform well under standard conditions, such as uniform backgrounds and consistent lighting (e.g., scanned documents). However, these techniques may fall short when dealing with text-in-wild scenarios [128], where the background and lighting are highly diverse and the text orientation is arbitrary.
For tasks involving the diverse image properties of natural scenes, such as identifying signs in urban areas [129], two primary subtasks are distinguished. Scene-text detection focuses on identifying and localizing text regions within an image without interpreting their content, while scene-text recognition converts these localized regions into machine-readable characters [130, 131, 132]. The combination of these tasks is known as scene-text spotting, which handles text detection and recognition in an end-to-end manner. These methods are generally faster and more efficient due to their unified feature extraction module [133, 134].
Focusing solely on character detection and recognition can be insufficient when image conditions are highly diverse, such as varying lighting and backgrounds. Relying on character detection in diverse scenes can increase false positives, as many patterns may look similar to letters. Additionally, DL methods require vast amounts of data, and character-level annotation is more expensive than word-level or line-level annotation due to its granularity.
Table 2: Main approaches to ILU identification. The table reports how often each combination appears in the surveyed literature. $\bullet$ = detection, $\bullet$ = recognition.
Object Character Text Count Percentage
20 54.1%
4 10.8%
3 8.1%
2 5.4%
2 5.4%
2 5.4%
1 2.7%
1 2.7%
1 2.7%
1 2.7%
This shift from character detection and recognition to text detection and recognition is shown in Table 2, where the majority of articles employ DL methods for both scene-text detection and recognition tasks. The share of hybrid approaches using non-DL methods for one of these steps are minimal. Examples include using traditional techniques for text detection such as cropping a predefined area for searching ID codes [127] or for character recognition, template matching after a DL model detects text areas [135]. Instead, researchers largely favour more comprehensive pipelines that rely on DL in each step. This shift towards unified DL-based approaches for both scene-text detection and recognition highlights their potential. However, to further enhance efficiency, researchers should focus more on text spotting [136, 134].
In ILU identification, text recognition remains the greatest challenge. Environmental conditions and technical issues such as unreliable hardware or connectivity problems between devices can cause delays at the gate. These delays increase waiting times for stakeholders, lead to congestion, add to environmental pollution, and create further setbacks for subsequent loading units. When recognition errors occur, the risk of incorrect inventory lists rises, which not only requires time-consuming manual corrections in the handling process but can also raise security concerns [14]. To address these challenges more effectively, end-to-end approaches have been proposed. By combining detection and recognition into a single forward pass with a unified feature extraction module, they streamline the process and improve overall efficiency.
Few studies introduce object detection as a prior step to scene-text detection and recognition [73, 80]. Detecting the ILU first reduces the search area for the text detection step and enables association of detected and recognised IDs with other objects, thereby improving scene understanding [73].
4.3 Datasets
Datasets form the backbone of reliable ILU detection. The quality, diversity, and annotation of the underlying data strongly influence how well models perform once deployed in real-world port environments. Carefully curated datasets with consistent labelling not only enable robust training and evaluation, but also determine whether methods can generalise beyond laboratory settings.
Understanding the properties of datasets is therefore essential for evaluating the performance and applicability of CV-based ILU identification systems. Below, we summarise the main characteristics and settings of the datasets used in the eligible studies, highlighting camera positioning, diversity, and availability.
4.3.1 Image Acquisition Setups
We categorised cameras as fixed or mobile cameras based on their mounting and usage. Fixed cameras include OCR gates, CCTV cameras, and crane-mounted cameras. These are installed at permanent locations such as terminal gates, yards, or crane structures, providing stable, controlled views for capturing ILU codes. Their main advantage is delivering consistent, high-quality images ideal for computationally efficient methods like classical DIP and traditional ML, as long as ILUs are present at predefined areas. Although the installation and operation of these systems involve substantial infrastructure investment, fixed cameras often struggle to capture sufficient feature information in large and complex environments, resulting in reduced accuracy and reliability of detection and recognition based on visual neural networks [137].
Figure 2: Sample images for fixed perspective (left) [72], ground vehicle perspective (middle) [73], aerial perspective (right) [84].
In contrast, mobile cameras are mounted on unmanned aerial vehicles, reach stackers, terminal trucks, or other industrial vehicles. For ground-level perspectives, we also considered handheld images of ILUs, as they closely resemble those captured by vehicle-mounted systems. Unlike fixed setups, mobile cameras provide flexibility by moving across the terminal and capturing a wider range of perspectives. This mobility enables real-time port monitoring and localisation of each ILU in real-world coordinates, as well as monitoring their positions by handling vehicles such as reach stackers [138, 139]. While they cover wider spatial areas and provide richer scene diversity, mobile cameras introduce more complex image variability, requiring advanced processing methods and diverse publicly available datasets.
Table 3: Summary of dataset characteristics: camera position and dataset availability.
Category Count Percentage
Camera position Only Fixed 47 74.60 %
Fixed and Ground 10 15.87 %
Fixed, Aerial and Ground 3 4.76 %
Only Ground 1 1.58 %
Only Aerial 1 1.58 %
Not reported 1 1.58 %
Availability Not public 54 85.71 %
Fully available 3 4.76 %
Partially available 3 4.76 %
Upon request 3 4.76 %
Behind paywall 1 1.58 %
As shown in Table 3 the most common collection setting is the fixed CCTV cameras at OCR-gates [140, 123, 141] or cranes (74.60 %) [80, 108] and a mixture of these images with mobile phone snapshots (15.87 %) [67]. However, vehicle-mounted cameras are increasingly used alongside fixed cameras to capture wider areas and provide information across the terminal [142, 84, 122, 73].
4.3.2 Diversity
Most datasets consist of images gathered from the real-world settings without controlled environments such as the World’s busiest ports like Shanghai [16], Singapore [44], Hong Kong [143], Kaohsiung [13], and Hamburg [73], as well as smaller inland ports like Dresden and Riesa [84]. Besides the terminal-specific data collection, web scraping were also used [125, 72]. In addition, controlled environments such as the Shanghai Maritime University test facility [42] were used for data collection.
Other datasets were created with images taken in Malaysia [66], Sweden [123], Vietnam [18, 124, 127, 68], Thailand [20], South Korea [142, 71, 144, 108, 102], Japan [99], and Italy [145]. However, datasets consisting of images from multiple locations are rare. For DL methods, capturing the nuances of ILUs across different countries and their respective ports can be important.
The number of images used in the experiments of the eligible articles ranges from a minimum of 30 images to a maximum of 34000, with a median of 1050 images. Since the DIP-based methods does not require a huge amount of data, researchers who tackled this problem before DL, experimented with a lower number of images compared to recent studies. However, with the emergence of ML and DL, the need for data has increased significantly. Most of the studies have addressed this need by pre-training their models on general-purpose scene-text detection and recognition datasets such as ICDAR2015 [146], CTW1500 [147], and SynthText [148], then fine-tuning on smaller ILU datasets.
Early datasets already included efforts to represent diverse conditions in order to evaluate generalisation. Lighting variability is the most discussed factor, with datasets often including images taken in bright, low-light, and uneven lighting scenarios [100, 43]. Considering how the earlier, thresholding-based methods may get affected by small changes in light intensity, diversifying the lighting conditions in the experiments was a major focus [145, 97, 143, 96]. However, currently lighting conditions are classified with more general classes: Day and night settings are frequently included to test the adaptability of the systems [104, 122, 73, 16].
Weather conditions such as rain, snow, and fog are also represented, as they can significantly affect image quality and ILU identification performance [73, 67, 126, 143]. Additionally, some data sets feature images with motion blur, camera noise, dirty, rusty, damaged, or partially occluded ID codes to further challenge algorithms with the exceptional conditions that exist in real-world settings [104, 73, 105, 126, 72]. Rarely were these conditions simulated to increase the data sets [125].
4.3.3 Availability
Table 3 shows the availability of the datasets used by the identified articles. The significant majority of the datasets are not publicly available, which creates a substantial barrier for research and development in this area. The lack of availability also prevents the reproducibility of results and fair comparison of the proposed methods. Further details for the publicly available datasets are shown in Table 4.
Among these, the TRansportation Unit Detection and Identification (TRUDI) dataset [73] stands out for its annotation density. It offers 35,034 instances across 733 images, covering five classes: Container, trailer, tank container, ID code, and logo. The Shipping Container Code dataset [23] leads in the number of images, and the Container Number-OCR dataset [105], though smaller (2,851 images from Tianjin Port), offers similar challenging conditions as others: variable weather conditions and degraded ID codes (dirt, rust, damage). Uniquely, TRUDI introduces motion blur as an additional challenging condition, as most of its images were captured via UAVs, reach stackers, and terminal trucks.
Table 4: Publicly available ILU identification datasets.
Dataset Images Instances Conditions Location
Shipping container code [23] 3600 4000 Day/night, Rainy, snowy Dirty, damaged ID Not reported
Container Number-OCR [105] 2851 2851 Low/high light Foggy Dirty, damaged ID Tianjin Port
TRUDI [73] 733 35034 Day/night Rainy, snowy, foggy Motion blur Dirty, damaged ID
Multiple countries,
Multiple terminals
4.4 Evaluation
To evaluate the proposed ILU identification methods, several evaluation metrics are commonly used. As there are multiple stages of ILU identification using CV methods, such as text detection and text recognition, each stage can be evaluated individually. However, end-to-end accuracy or success rate, is the most relevant metric due to its practical importance. It is the most strict metric as one misrecognised character is enough for failed identification.
Since the ILU identification problem is mostly solved with multiple stages such as text detection and text recognition, some articles only report the results of individual stages and do not include end-to-end results. However, the results for individual stages offer limited insight for the applicability of the method for the ILU identification task.
Precision and recall are the common metrics for these individual stages. These metrics show the balance between false positives and false negatives, with precision focusing on the accuracy of positive predictions and recall on the model’s ability to identify all relevant instances. The F1-score, or H-mean, harmoniously combines precision and recall into a single metric, offering a balanced view of the model’s performance.
For object detection or text detection tasks, mean average precision (mAP) and average precision (AP) are commonly used, measuring precision across various intersection over union (IoU) levels. The IoU metric is important for assessing the overlap between predicted and actual bounding boxes. As for the recognition, besides precision, recall and F1-score, other metrics such as the character error rate (CER) is used. CER measures the proportion of characters that were incorrectly predicted compared to the total number of characters in the reference text, using the Levenshtein-distance [149].
Speed and efficiency are evaluated using metrics such as frames per second (FPS), and inference time, which assess the model’s operational speed. Model complexity is assessed through the number of parameters, giga-floating point operations per seconds, and model size, indicating the computational demands and storage requirements.
Although the recent studies use the same, established metrics adapted from general purpose text detection and recognition task, putting the reported results for ILU identification into a frame is not possible with the limited amount of publicly available datasets. Due to different datasets used in the experiments, the reported end-to-end accuracy of the proposed methods ranges from 5 % [142] to 96,17 % [144]. This level of variance can be observed in every metric, including the processing speed.
4.5 Future Research Directions
In the literature, future directions for the ILU identification task mainly focus on enhancing the datasets by including more images taken from diverse environments such as various distances, angles, and under different weather and lighting conditions. This expansion aims to cover multiple ports and scenarios to improve model generalisation. However, increasing the diversity of proprietary datasets would not solve the fundamental problem with the datasets for ILU identification. Recent papers acknowledge the lack of benchmark datasets and its negative effects on the field [122, 73]. Given that creating a new dataset for downstream CV tasks are tedious, time consuming, and costly [150, 151], alternatives such as synthetic data generation can be considered [152, 148], e.g., by using a game engine [153].
Handling challenging conditions, such as low-light scenarios, occlusions, and damaged ID codes, is another direction pointed by the researchers, with techniques proposed to manage partial occlusions and degraded text [105]. Current solutions are relying on ”off-the-shelf” super-resolution methods such as Super-Resolution-CNN [154]. Literature on scene-text super-resolution has focused primarily on natural language, leaving an opportunity for research specifically on contextless text such as ILU codes. Alternatively, multi-view recognition is also identified as a future direction due to its potential against degraded ID codes [23, 12] as multiple images of a single ILU are captured from different angles in this approach.
Improving image quality through preprocessing is another focus area. Suggested methods include image enhancement techniques such as noise reduction [17], or the use of Generative Adversarial Networks [155, 156] to address issues such as shadows, rust, and dirt [23, 105]. Postprocessing steps are also recommended to verify recognition results by using the check digit and the owner code. [15, 157, 69, 72].
For DL-based methods, refining network architectures to increase the receptive field so that the model can see more of the scene in a single pass. It is important for detecting containers appears to be small due to perspective [122]. Another future work suggestion is incorporating attention mechanisms to boost recognition accuracy [158]. The development of unified, end-to-end models is suggested to streamline processing pipelines [73, 140, 105, 157].
The contextless nature of ISO6346 codes makes scene-text spotting task harder as there are no natural language cues [159]. Therefore, directly applying current scene-text recognition models that rely on semantic context, is not the most suitable approach, as they usually use a language model or at least a decoder module pretrained on natural language to correct the initial recognition result [160, 83]. However, incorporating complex language models increases model complexity, while not yielding any benefits for contextless recognition tasks [159]. Researching on the contextless scene-text recognition models is an opportunity which may result in novel and more efficient methodologies. These new methods can also be useful for similar applications including International Union of Railways (UIC)-wagon codes, or registration plates.
Improvements for real-time processing are also highlighted, with a focus on optimising algorithms for mobile devices and enhancing systems to work in real-time, potentially through video stream analysis [123, 42].
5 Conclusion
This review provides the first synthesis of CV-based ILU identification research by analysing 63 empirical studies published over three decades. There is a clear upward trend for the number of publications with the rise of DL methods and Asia’s dominance as a global trade hub, with 79.71 % of studies originating from the region. Public funding supports over a third of research highlighting a reliance on public investment for transportation related infrastructure. Policymakers could foster progress by funding collaborative initiatives and public-private partnerships, which currently account for just 6.35 % of studies.
Although there are articles published in top journals and conferences, most studies were published in unranked outlets. The main reason for this is the direct application of general purpose CV methods to the ILU identification task. This lack of methodological contribution can be increased by adapting existing methods with modifications that addresses challenges, such as contextless text spotting and real-time processing constraints.
Methodologically, the field has shifted from classical image processing to DL, currently used in more than 90 % of recent studies, thanks to GPU advancements and frameworks like PyTorch. This shift comes with the need of huge datasets to train DL models. Data limitations are the biggest obstacle with 85.71 % of datasets being private and most studies relying on controlled image acquisition setups. The median dataset size is too small for robust DL training, forcing researchers to adapt general text detection and recognition models. Another critical point for improving the quality and credibility of publications is using or developing benchmark datasets. The lack of benchmark dataset makes performance comparisons difficult, leaving gaps in understanding which methods work best in which setting.
The research on ILU identification mainly focuses on fixed camera settings such as OCR-gates. However, recent studies show a shift from fixed to mobile cameras, such as drones or vehicle mounted cameras. This shift can spark more innovation due to new challenges related to dynamic scenes from the technical perspective [84] as well as social perspective such as privacy issues [161, 162, 163]. Camera mounted vehicles offer new subtasks for the ILU identification task such as pose estimation of the ILUs by predicting the real-world coordinates (position) and heading (orientation). This way, a precise monitoring of the terminal is possible.
To conclude, we suggest emphasising standardisation for evaluation with publicly available benchmark datasets, contextless text spotting architectures, real-time processing on edge devices and more focus on vehicle mounted camera setups for a more comprehensive ILU identification and terminal monitoring.
Acknowledgements
The project is supported by the German Federal Ministry for Digital and Transport (BMDV) in the funding program Innovative Hafentechnologien II (IHATEC II) with the funding number 19H23002C.
Disclosure statement
The authors report there are no competing interests to declare.

### v2

@@
 Abstract
-The standardisation of Intermodal Loading Units (ILUs), such as containers, semi-trailers and swap bodies, has revolutionised global trade yet their efficient and robust identification remains a critical bottleneck in high-throughput ports and terminals.
-This paper reviews 63 empirical studies that propose computer vision (CV) based solutions.
-It covers the last 35 years (1990-2025), tracing the field’s evolution from early digital image processing (DIP) and traditional machine learning (ML) to the current dominance of deep learning (DL) techniques.
-While CV offers cost-effective alternatives for other types of identification techniques, its development is hindered by the lack of publicly available benchmarking datasets.
-This results in high variance for the reported results such as end-to-end accuracy ranging from 5 % to 96 %.
-Beyond dataset limitations, this review highlights the emerging challenges especially introduced by the shift from character-based text recognition to scene-text spotting and the integration of mobile cameras (e.g.
-drones, sensor equipped ground vehicles) for dynamic terminal monitoring.
-To advance the field, the paper calls for standardised terminology, open-access datasets, shared source code, while outlining future research directions such as contextless text recognition optimised for ISO6346 codes.
+Background: The standardisation of Intermodal Loading Units (ILUs), including containers, semi-trailers, and swap bodies, has transformed global trade, yet efficient and robust identification remains an operational bottleneck in ports and terminals.
+Objective: To map Computer Vision (CV) methods for ILU identification, clarify terminology, summarise the evolution of proposed approaches, and highlight research gaps, future directions and their potential effects on terminal operations.
+Methods: Following PRISMA-ScR, we searched Google Scholar and dblp for English-language studies with quantitative results.
+After dual reviewer screening, the studies were charted across methods, datasets, and evaluation metrics.
+Results: 63 empirical studies on CV-based solutions for the ILU identification task, published between 1990 and 2025 were reviewed.
+Methodological evolution of ILU identification solutions, datasets, evaluation of the proposed methods and future research directions are summarised.
+A shift from static (e.g.
+OCR-gates) to vehicle mounted camera setups, which enables precise monitoring is observed.
+The reported results for end-to-end accuracy range from 5 % to 96 %.
+Conclusions: We propose standardised terminology, advocate for open-access datasets, codebases and model weights to enable fair evaluation and define future work directions.
+The shift from static to dynamic camera settings introduces new challenges that have transformative potential for transportation and logistics.
+However, the lack of public benchmark datasets, open-access code, and standardised terminology hinders the advancements in this field.
+As for the future work, we suggest addressing the new challenges emerged from vehicle mounted cameras, exploring synthetic data generation, refining the multi-stage methods into unified end-to-end models to reduce complexity, and focusing on contextless text recognition.
+Addressing these challenges can result in minimised human errors, decreased turnaround times, reduced operational costs and higher resilience against disruptions.
 keywords:
-intermodal loading unit identification , container code recognition , text detection , text recognition , logistics automation
-journal: Transport Reviews
+intermodal loading unit identification , container code recognition ,
+journal: Transportation Research Part: C
 ILU
@@
 unmanned aerial vehicle
-IGD
-InteGreatDrones
 CER
@@
 giga-floating point operations per second
-ID
-identifier
 TRUDI
@@
 Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews
-CMT
-Containerised Maritime Transport
 PCC
@@
 Deep Learning
-AI
-Artificial Intelligence
 CV
@@
 Automatic Container Code Recognition
-ALPR
-Automatic Licence Plate Recognition
 SVM
@@
 Generative Adversarial Network
-CCA
-Connected Component Analysis
-MSER
-Maximally Stable Extremal Regions
 SIFT
@@
 Histogram of Gradients
-LBP
-Local Binary Pattern
 […]
@@
 1 Introduction
-The evolution of cargo transportation has been significantly transformed by the concept of containerisation.
-Before, cargo handling was labour-intensive, neither standardised nor suitable for automatisation.
-Handling and storage arrangements were costly and required special attention [1, 2], resulting in terminal congestion.
-Containerisation has facilitated automation and improved operational efficiency in cargo handling [3], while also significantly contributed to the growth of global trade [4].
-The standardisation of freight handling has not only led to the shipping container but also led to a variety of standardised types of Intermodal Loading Units.
-The term ILU has been introduced to describe units designed to be easily transferred between different modes of transport, such as vessels, trains, and trucks, without the need to unload and reload the contents.
-The loading units, such as containers, interchangeable bodies (swap bodies) and semi-trailers, fall under this term [5].
-These units are typically built according to specifications defined in standards such as ISO668 [6].
-These specifications ensure compatibility of ILUs with different transport systems [7] for smooth movement of goods along supply chains [8].
+The term Intermodal Loading Unit (ILU) has been introduced to describe units designed to be easily transferred between different modes of transport, such as vessels, trains, and trucks, without the need to unload and reload the contents.
+The loading units, such as containers, interchangeable bodies (swap bodies) and cranable semi-trailers, fall under this term [1].
 A unique identification code, following the ISO6346 [2] standard, is assigned to each ILU for tracking purposes.
-This code consists of four letters, a six-digit number and a check digit, which together indicate the owner and type of the ILU.
+This code consists of four letters, a six-digit number and a check digit, which together indicate the owner, type and serial number of ILU.
+Identically looking ILUs can exclusively be distinguished by their ID codes.
+While the standardised identification of ILUs provides a framework for tracking, the scale of modern port operations introduces challenges.
 As global trade expands, many leading ports handle tens of millions of containers annually, with some operating at or above their nominal design capacity [3].
 In this high throughput, keeping track of ILUs is important to ensure efficient management [4].
-The often identically looking ILUs can exclusively be distinguished by their ID codes.
-Therefore, a reliable system that can rapidly recognise these codes is indispensable.
+For example, at transfer points of the terminal to other modes of transport, each ILU must be identified and cross-checked with its associated documentation before entry or exit.
+When automated systems fail to achieve reliable ILU identification, manual verification becomes necessary, further slowing down the process.
+Even a brief slowdown at a busy gate can cause truck queues stretching hundreds of metres, disrupting yard operations and increasing turnaround times for transport vehicles [5, 6].
+Extended waiting times contribute to congestion which increases fuel consumption and greenhouse gas emissions from idling trucks, amplifying the environmental footprint of port activities [7].
+Therefore, efficient and automatic ILU identification is essential for overcoming these bottlenecks and sustaining the growth of containerised transport.
 Standardisation and digitalisation enabled the development of various techniques for ILU identification.
 Among these, Computer Vision (CV)-based identification code recognition stands out as a cost-effective alternative to RFID, which achieves only about 70 % accuracy and entails high installation and maintenance costs [8, 9].
-In contrast, CV-based methods can be implemented using a range of camera setups from high-end static Optical Character Recognition (OCR) gates [14], that can identify entering and exiting ILUs [15, 16], to more affordable and flexible options such as using the existing CCTV infrastructure, cameras mounted on aerial or ground vehicles, or even mobile devices, which significantly reduce the need for specialised hardware.
-Furthermore, CV-based identification plays a crucial role in enhancing security by providing accurate verification for customs clearance, which can support in the correct declaration of goods [17, 18, 19].
-Although CV-based methods reduce human-related errors and increase the processing speed [20, 21, 22, 23] there are several challenges: Containers often arrive in rapid succession, will be stacked for storage, and can be positioned at angles that partially occlude their ID codes.
-Environmental factors such as rain, fog, poor lighting, or glare from sunlight can degrade image quality, while harsh conditions, such as dirt, rust, or physical damage, further complicate automated ID code recognition [14, 24].
-In addition, variations in font style, colour and placement of ID codes across different operators and regions create inconsistencies that reduce the reliability of automatic recognition systems.
-At transfer points of the terminal to other modes of transport, each ILU must be identified and cross-checked with its associated documentation before entry or exit.
-When automated systems fail to achieve reliable recognition, manual verification becomes necessary, further slowing down the process.
-Even a brief slowdown at a busy gate can cause truck queues stretching hundreds of meters [25], disrupting yard operations and increasing turnaround times for transport vehicles.
-Extended waiting times contribute to congestion which increases fuel consumption and greenhouse gas emissions from idling trucks, amplifying the environmental footprint of port activities [26].
-Therefore, efficient ILU identification is essential for overcoming these operational bottlenecks and sustaining the growth of containerised transport.
-Despite its crucial role in improving transportation and trade security, to the best of our knowledge, there are not any literature reviews on ILU identification.
+In contrast, CV-based methods can be implemented using a range of camera setups as shown in Figure 1.
+From static Optical Character Recognition (OCR) gates [10, 11], that can identify entering and exiting ILUs [12, 13], the field is shifting towards more flexible options such as camera equipped aerial or ground vehicles, or even mobile devices, which significantly reduce the need for specialised infrastructure [14, 15, 16, 17].
+Furthermore, CV-based ILU identification plays a crucial role in enhancing security by providing accurate verification for customs clearance, which can support in the correct declaration of goods [18, 19, 20].
+Figure 1: Schematic overview of a terminal, highlighting the use of camera-equipped aerial and ground vehicles and OCR-gates for comprehensive port monitoring.
+Despite its crucial role in improving port operations by reducing human-related errors and increase the processing speed [21, 22, 23, 24], to the best of our knowledge, there are not any literature reviews on CV-based ILU identification.
 To fill this gap, we prepared this literature review by following Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) guidelines [25] with the intention of identifying the available publications and proposed concepts.
@@
   • 1.
-    Catalogue all empirical studies on Intermodal Loading Unit (ILU) identification using Computer Vision (CV) methods between 1990 and 2025.
+    Catalogue all empirical studies on ILU identification using CV methods between 1990 and 2025.
   • 2.
@@
   • 3.
-    Synthesise an overview of Intermodal Loading Unit (ILU) identification methods, image acquisition settings, datasets, and evaluation metrics used in these studies.
+    Synthesise an overview of ILU identification methods, image acquisition settings, datasets, and evaluation metrics used in these studies.
   • 4.
@@
 2 Terminology
-This review distinguishes the roles of Digital Image Processing (DIP), Machine Learning (ML), and Deep Learning (DL) in the context of CV-based ILU identification to ensure precise use of technical terms and prevent misinterpretation across disciplines.
-Therefore, we provide the definitions of the general approaches:
+As CV-based ILU identification is an interdisciplinary field we would like to establish a common terminology to create a common ground for researchers with different background, such as robotics, artificial intelligence, and logistics.
+In the related literature, we realised a lack of definitions which results in ambiguity.
+For example, text spotting, an integrated task encompassing both text detection and text recognition, is commonly reduced to OCR, ignoring its broader scope.
+This ambiguous terminology complicates result interpretation, and hindering meaningful cross-study comparisons.
+Therefore, we provide the precise definitions for general approaches and downstream tasks.
+2.1 Approaches
 Digital Image Processing (DIP): Classical, rule-based image operations (e.g., denoising [26], contrast normalisation [27], edge detection [28, 29], morphologic operations [30], thresholding [31, 32], geometric transforms [33]).
@@
 non-object, scene categories).
-Common models include Support Vector Machines [45, 46, 47], Random Forests [48], decision trees [49] and k-Nearest Neighbors [50].
-These reduce manual rule-tuning and can generalize better under variations in viewpoint, scale, and lighting, but still depend critically on quality of the features and consistency in preprocessing [51].
-Deep Learning (DL): A subfield of ML using deep neural networks, such as Convolutional Neural Networks [52] or transformers [53] that learn features and classifiers jointly and end-to-end.
-Although DL requires substantial training data and computational resources it automatise the feature engineering step.
+Common models include Support Vector Machines [43], Random Forests [44], decision trees [45], k-Nearest Neighbors [46] and Reinforcement Learning [47, 48, 49].
+These reduce manual rule-tuning and can generalise better under variations in viewpoint, scale, and lighting, but still depend critically on quality of the features and consistency in preprocessing.
+Deep Learning (DL): A subfield of ML using deep neural networks, such as Convolutional Neural Networks [50, 51] or Transformers [52] that learn features and classifiers jointly and end-to-end.
+Although DL requires substantial training data and computational resources, it automates the feature engineering process.
+However, a major drawback of deep learning models is their black-box nature, which makes them difficult to interpret and explain [53].
 For similar tasks depending on similar features, deep learning methods can benefit from transfer learning [54] and fine-tuning [55].
-Besides the terminology for general approaches, we would like to provide definitions for CV-specific terminology:
+2.2 ILU Identification Related Computer Vision Tasks
+Besides the terminology for general approaches, we would like to provide definitions for downstream tasks:
 Object detection: Localising (drawing bounding boxes around) and classifying (assigning labels to) objects within an image.
@@
 Similar to the object detection, character detection provides information about where individual characters are and which characters they are [61].
-(Scene-)Text detection: Localising regions containing text in an image (e.g., street signs,) without extracting its textual content.
-Outputs bounding boxes (often quadrilaterals) around words or lines.
-Unlike character detection, it focuses on full words or lines, which makes it more robust against distractors [62].
-(Scene-)Text recognition: Extracting textual information from detected text regions by converting them into machine-readable characters or words [63, 64].
+Optical Character Recognition (OCR): Localising and identifying individual characters in an image.
+OCR is predominantly used for document text recognition and perform well under standard conditions, such as uniform backgrounds and consistent lighting (e.g., scanned documents).
+However, OCR techniques may fall short when dealing with text-in-wild scenarios [62], where background and lighting are highly diverse and text orientation is arbitrary.
+Relying on character detection in diverse scenes can increase false positives, as many objects share similar patterns with individual letters [63, 64].
+Another disadvantage of character-level recognition, especially with DL methods, is that character-level recognition needs more expensive annotated data than word-level or line-level recognition due to its granularity and required labour for annotating individual characters.
+For tasks involving the diverse scenes, such as identifying signs in urban areas [65], two primary subtasks are defined:
+(Scene-)Text detection: Localising regions containing text in an image without extracting its textual content.
+Outputs quadrilaterals around words or lines.
+Unlike character detection, it focuses on full words or lines, which makes it more robust against distractors [66, 67, 68, 69].
+(Scene-)Text recognition: Extracting textual information from detected text regions by converting them into machine-readable characters or words [70, 71, 67, 68, 69, 69].
 Text spotting: End-to-end method that jointly detects and recognises text in a single step [72, 61].
 2.3 Terminology Beyond OCR Gates
-Initially, CV methods were applied at terminal gates, where the task is to automatically identify loading units as they entered or exited the facility.
-These gates are among the most vulnerable and critical points of a terminal, making reliable identification essential.
-The task performed there was to capture an image of the loading unit and read its identification code through DIP [44].
-In the literature, this task is often referred to as Automatic Container Code Recognition (ACCR) [66, 67, 68, 69, 70, 71, 72] or transportation unit (TU) identification [73].
-The focus of ACCR is on using CV to read the ISO6346 identification codes from containers at the point of entry.
-These entrances remained as the critical bottlenecks in terms of operational efficiency and congestion impact, so their optimisation becomes one of the key challenges in terminal management [74].
-However, the term ACCR is not precise.
-First, the word container is too narrow, since not only containers but also semi-trailers and swap bodies carry ISO6346-compliant codes (as defined in EN 13044 [75]).
+Terminal entrances remained as the critical bottlenecks in terms of operational efficiency and congestion impact, so their optimisation becomes one of the key challenges in terminal management.
+The earlier studies on this, applied CV methods only at terminal gates, where the task is to automatically identify loading units as they entered or exited the facility.
+In the literature, this task is referred to as Automatic Container Code Recognition (ACCR) [73, 74, 75, 76, 77, 78, 79] and rarely as loading unit identification or transportation unit (TU) identification [17].
+The focus of ACCR is on containers at terminal gates.
+However, the word container does not cover all types of loading units such as semi-trailers and swap bodies that has ISO6346-compliant codes (as defined in EN 13044 [80]).
 Second, the word recognition does not fully capture the actual tasks involved.
-In computer vision, recognition usually means extracting text from a clean image, especially for the mobile cameras.
 The term recognition refers to extraction of text from an image in the CV field.
@@
 Over time, these methods were extended beyond the gates.
-With advancements in computer vision and robotics, identification is now performed at quay cranes [77], yard blocks [78], and movable objects such as reach stackers [79].
-These new use cases introduce more variation and complexity than the original gate-based setting.
-For these reasons, we propose the term Intermodal Loading Unit (ILU) identification as an inclusive designation for the task of determining which specific ILUs with ISO 6346–compliant ID codes are visible in an image.
-This term includes various loading units and also for the CV terminology such as ILU detection/segmentation stage introduced in recent papers [73, 80], scene-text detection and recognition.
+With advancements in CV and robotics, identification is now performed at quay cranes [82], yard blocks, and terminal vehicles such as reach stackers [83].
+These new use cases introduce more variation and complexity than the original gate-based setting and therefore, text recognition is not sufficient by itself.
+For these reasons, we propose the term Intermodal Loading Unit Identification as an inclusive designation for the task of determining ID codes of ILUs.
+This term covers various loading units and also ILU detection/segmentation stage prior to text detection and recognition steps, introduced in some studies [17, 84].
 A term that covers all these stages aligned with CV terminology is important, as it would address shared challenges across the CV applications.
-For example, using a common terminology can link the relevant studies to the broader problems such as small object segmentation [81] or small text detection [82] which are actively studied in general purpose CV methods.
-Moreover, a common terminology can help in finding the distinct aspects of ILU identification.
-Unlike in most common text recognition needs, where a language model based encoder (e.g.
-TrOCR [83]) is useful due to contextual cues present in the natural language, the contextless nature of the target text in ILU identification makes this approach ineffective.
+Using a common terminology can link the relevant studies to the broader problems.
+For example, different scales of objects and text is a problem that was introduced with the mobile cameras for ILU identification task.
+However, small object segmentation [85] or small text detection [86] havebeen actively studied in general purpose CV methods.
+Moreover, a common terminology can help in finding the distinct aspects of ILU identification from the broader tasks.
+Unlike in common text recognition, where a language model based encoder (e.g.
+TrOCR [87]) is useful due to contextual cues present in the natural language, the contextless nature of ISO6346-compliant ID codes makes this approach ineffective.
 Such models tend to overfit to frequently occurring text instances in training data [15], limiting their applicability in this context.
 Adopting the term ILU identification can provide conceptual clarity and establish a coherent basis for researchers and practitioners.
+Having established the terminology and challenges, we now turn to the methodology used for systematically review the literature.
 3 Methodology
+This literature review was conducted in accordance with the PRISMA-ScR guidelines [25] and eligibility criteria were specified a priori and piloted before the screening began.
+These criteria follow the Population–Concept–Context (PCC) logic and Table 1 details the inclusion and exclusion rules together with their rationales.
+Table 1: Inclusion and exclusion criteria for systematic review of ILU code recognition.
+Domain Inclusion rule Exclusion rule
+Date range 1 Jan 1990–31 August 2025 << 1990 or after cut-off
+Language English full text Non-English; abstract-only
+Publication type Peer-reviewed journal / conference paper; pre-print (quantitative) Patents, standard, theses, news or editorials
+Population (P) ILUs (containers, swap-bodies, semi-trailers) Other objects unless ILUs also analysed
+Concept (C) CV localisation and recognition of alphanumeric ILU code RFID/barcode solutions
+Context (C) Fixed or mobile cameras in ports, terminals, rail yards Depots
+Reporting completeness Method detail plus numeric outcomes (precision, recall, F1, mAP, read-rate) No quantitative results
+Duplicates / accessibility First complete, accessible full text Duplicates; inaccessible full text
+As the first review on ILU identification, we applied inclusive eligibility criteria to capture the full scope of available evidence.
+Our review includes articles written in English, published from 1990 to 2025, that follow a scientific structure.
+Both peer-reviewed articles and preprints are included as long as they report quantitative results, allowing us to catalogue the latest research findings.
+The final selection of articles and data charting were determined through consensus among the authors, following a review of titles, abstracts, and full texts.
 We finalised our search for articles in August 2025, using Google Scholar and dblp-Computer Science Bibliography to find scientific articles on ILU identification using CV or DIP techniques.
@@
 We checked the cited papers and ’related articles’ in Google Scholar and Research Rabbit to expand our pool of eligible articles.
-As the first review on ILU identification, we applied inclusive eligibility criteria to capture the full scope of available evidence.
-These criteria were specified a priori before screening began.
-They follow the Population–Concept–Context (PCC) logic and the methodological guidance in [27].
-Our review includes articles written in English, published from 1990 to 2025, that follow a scientific structure.
-Both peer-reviewed articles and preprints are included as long as they report quantitative results, allowing us to catalogue the latest research findings.
-The final selection of articles and data charting were determined through consensus among the authors, following a review of titles, abstracts, and full texts.
-From the 63 eligible articles the following data items were extracted into a spreadsheet [^1] https://cloud.uni-hamburg.de/s/FHqaGLR4Ccy8ceG: journal or conference details, contributions, methods used, method class (DIP, ML, or DL), dataset specifics (availability, number of images, diversity in lighting and weather), image collection region and camera position such as fixed (CCTV) or moving (aerial or ground vehicle), evaluation metrics, results, deployment status, suggested future directions and funding.
+From the 63 eligible articles the following data items were extracted into a spreadsheet [^1] Available in supplementary materials.: journal or conference details, contributions, methods used, method class (DIP, ML, or DL), dataset specifics (availability, number of images, diversity in lighting and weather), image collection region and camera position such as fixed (CCTV) or mobile (aerial or ground vehicle), evaluation metrics, results, deployment status, suggested future directions and funding.
 The resulting table was then analysed and discussed by the authors.
@@
 During the screening process, articles were excluded due to being duplicates ($n=2$), not reporting quantitative results ($n=6$), written in a non-English language ($n=5$), focusing solely on container detection or integrity assessment rather than identification via ISO6346 compliant codes ($n=6$), solutions for other industrial text recognition applications such as wagon code recognition ($n=5$), unavailability of their full text ($n=2$), or being patents ($n=2$).
-Table 1: Eligible article characteristics ($n=63$). Five articles involve international collaborations, which results in the count of 69 rather than the 63, in geographic regions section.
-Category Characteristic Count Percentage
-Year of publication 1990–1995 2 3.17 %
-1996–2000 1 1.59 %
-2001–2005 5 7.94 %
-2006–2010 3 4.76 %
-2011–2015 10 15.87 %
-2016–2020 18 28.57 %
-2021–2025 24 38.10 %
-Geographic region Asia 55 79.71 %
-Europe 8 11.59 %
-North America 3 4.35 %
-Africa 2 2.90%
-Oceania 1 1.45 %
-Funding source Not reported 32 50.79 %
-Public 24 38.10 %
-Public-Industry Collab. 4 6.35 %
-Industry 3 4.76 %
-Rank Unranked 32 50.79 %
-Q1 - Scimago 8 12.70 %
-A - CORE 5 7.94 %
-Q2 - Scimago 5 7.94 %
-B - CORE 4 6.35 %
-Q3 - Scimago 4 6.35 %
-Q4 - Scimago 4 6.35 %
-C - CORE 1 1.59 %
-Table 1 presents the characteristics of the eligible articles.
-The data indicate a consistent increase in publications over time, especially from 2011 onwards.
-This trend reflects growing interest in automatic ILU identification systems, likely driven by advancements in CV as well as greater computational power affordable to major terminals and ports.
-Asia contributes 79.71 % of the publications.
-This region handles high volumes of maritime traffic requiring efficient logistics and ILU identification systems.
-Countries such as China, Taiwan, South Korea, Vietnam, and Japan play significant roles in global trade, making effective logistics and supply chain management essential.
-Substantial investment in research and development by Asian governments and companies further supports this trend [85, 86].
-Public funding is reported as the main source, supporting about two out of five of the publications, indicating strong governmental backing.
-This support is often motivated by the potential for technological advancements to benefit national infrastructure and economic growth [87, 88, 89].
-Public funding is particularly important for the logistics sector that operates with low profit margins [90, 91].
-Industry funding and public-industry collaborations constitute one ninth of the publications, suggesting that while there is some industry involvement, companies largely rely on public sector investments to drive technological development.
-However, the presence of pure industry funding indicates a genuine market demand for companies to invest in research to remain competitive [92, 11].
+Figure 2: Geographic distribution of articles by country (n=63). Five articles involve international collaborations, which results in 69 for country mentions rather than 63.
+Most of the published articles were written by the authors with Asian affiliations (79,71 %) as shown in Figure 2.
+This is related with the high volumes of ILU handling in Asian ports and terminals, which creates a need for efficient ILU identification systems.
+Countries such as China, Taiwan, South Korea, Vietnam, and Japan play significant roles in global trade.
+This leads to innovation as reflected by the contribution of Asian countries to the scientific publications.
+Substantial investment in research and development by Asian governments and companies further supports this innovation [88, 89].
+Public funding is reported as the main source, supporting about two out of five publications, indicating strong governmental backing for the logistics sector that operates with low profit margins [90, 91].
+The motivation for strong governmental backing is the potential for technological advancements that will benefit national infrastructure and economic growth [92, 93].
+Besides this dominant public sector support, industry funding and public-industry collaborations constitute one ninth of the publications, suggesting that while there is some industry involvement, companies largely rely on public sector investments to drive technological development of ILU identification systems.
+Nevertheless, the presence of pure industry funding (4,76 %) indicates a genuine market demand for companies to invest in research to remain competitive [94, 4].
+Figure 3: Ranking of publication outlets for the eligible articles ($n=63$) over the year.
+The years with zero eligible publications are ommited in this barchart.
+As this scoping review was prepared in 2025, the data for that year remain incomplete, covering only publications up to August.
+While funding sources indicate industry and governmental priorities, the ranking of publication venues offers insight into the field’s academic recognition.
 To analyse the rankings of journals and conferences we followed the rankings of Scimago [^2] https://www.scimagojr.com/ for journals and CORE [^3] https://www.core.edu.au/icore-portal for conferences.
-Table 1 shows that more than half of the articles on automatic ILU identification task are published in unranked outlets, which is not uncommon for applied or niche technical topics [93].
-Among the ranked publications, Q1-Scimago and A-CORE have the highest numbers, suggesting that the topic is relevant for high-quality research venues.
-However, the majority of publications are spread across lower-tier or unranked venues.
-This distribution implies that the field could benefit from a greater emphasis on more novel contributions, increased comparability, and reproducibility of the results.
+Figure 3 shows that more than half of the articles on automatic ILU identification task are published in unranked outlets, which is not uncommon for applied or niche technical topics [95].
+Especially, the early years of DL dominance marked a period of stagnation (2017-2019), where the field struggled to deliver methodological innovation.
+This trend is changing in the recent years as the percentage of articles published in ranked outlets are increasing, as well as the total number of publications, suggesting that the ILU identification topic is relevant for high-quality research venues.
+Nevertheless, the field could benefit from a greater emphasis on more novel contributions, increased comparability, and reproducibility of the results.
 This could be achieved through publicly available benchmark datasets and repositories that showcase the source code and the model weights.
 By adopting such research standards, ILU identification tasks with CV-based methods could be more relevant for high-ranked publication outlets.
-Overall, the characteristics of the reviewed articles highlight a growing research field characterised by substantial activity in Asia and a predominance of public funding.
-Expanding funding sources and promoting industry collaboration may further advance automatic ILU identification through CV technologies.
 4.1 Methods
-Figure 1: Approach categories used for ILU identification task over the years.
 The development of methods for ILU identification has undergone a significant transformation over three decades, as illustrated in Figure 4.
@@
 These methods performed adequately in controlled image acquisition systems like OCR-gates [105] but struggled with challenging conditions, such as variability of brightness [75] and damages [106, 107] on the ILU images.
-In the subsequent five years, a modest increase in the use of ML can be observed, particularly in hybrid DIP and ML approaches [106, 107, 108, 109, 110, 12].
-This reflects the growing recognition of ML’s potential for enhancing robustness in character and symbol recognition tasks [111] in CV.
-However, the field lagged several years behind the broad adoption of DL for other applications, despite its transformative influence on CV [112, 113] from 2012 onwards (e.g., the emergence of AlexNet [114]).
-This delay may be attributed to the scarcity of labelled ILU identification datasets, limited computational infrastructure, and the conservative approach for adopting new technologies in the logistics industry.
+Figure 4: The number of publications and their approach categories used for ILU identification task over the years.
+In the subsequent five years, a modest increase in ML-based methods can be observed, particularly in hybrid DIP and ML approaches [108, 109, 110, 111, 112, 8].
+This reflects the growing recognition of ML’s potential for enhancing robustness in character and symbol recognition tasks in CV [113].
+However, the field lagged several years behind the broad adoption of DL, despite its transformative influence on CV solutions [114, 115] from 2012 onwards (e.g., the emergence of AlexNet [116]).
+This delay may be attributed to the scarcity of labelled ILU identification datasets, limited computational infrastructure, lack of technology skills in the workforce, and the conservative approach for adopting new technologies in the logistics industry [117, 118].
 The five-year interval from 2016 to 2020 marks a turning point, with a notable adaptation of DL-based methods.
@@
 As a consequence, purely DIP-based approaches declined significantly, often limited to pre- or post-processing roles within hybrid pipelines [75, 84, 21, 124, 77].
-ILU identification with fixed cameras often resembles OCR from scanned documents, as the background is mostly uniform.
-Segmenting and recognising characters is a sensible approach in such cases.
-Until 2016, all methods developed for ILU identification focused on character-level localisation and recognition, and this approach is still in use with DL methods [122, 123, 124].
-However, with the increasing use of mobile cameras (e.g.
-vehicle-mounted cameras), the task has evolved to resemble scene-text detection and recognition.
+Table 2: Stages defined in DL-based pipelines for ILU identification. The table reports how often each combination appears in the surveyed literature. $\bullet$ = detection, $\bullet$ = recognition.
+Object Character Text Count Percentage
+20 54,1%
+4 10,8%
+3 8,1%
+2 5,4%
+2 5,4%
+2 5,4%
+1 2,7%
+1 2,7%
+1 2,7%
+1 2,7%
+Until 2016, all methods developed for ILU identification focused on character-level localisation and recognition with fixed camera settings.
+This approach is still in use with DL methods [16, 125, 126] as ILU identification on fixed cameras often resembles OCR on scanned documents.
+However, with the increasing use of vehicle-mounted cameras, the task has evolved to resemble scene-text detection and recognition rather than OCR.
 As a result of this, the trend is shifting towards using comprehensive scene-text detection and recognition methods rather than relying on character detection and recognition.
-By 2021–2025, DL-based approaches had established themselves as the predominant paradigm for ILU identification.
-Even hybrid DIP-based methods virtually disappeared, indicating that the DL models now offer both the required accuracy and operational robustness [66, 125, 126].
-The relative infrequency of hybrid approaches suggests that most effective solutions now rarely require DIP or traditional ML methods, except in specialised scenarios such as real-time constraints [104, 16, 18, 127, 124, 15].
-4.2 Main Approaches for Deep Learning
-Observing the growing interest of DL methods, we now focus on the CV tasks relevant for extracting textual information.
-OCR techniques are designed to identify and localize individual characters in an image.
-They are used predominantly for document text recognition and perform well under standard conditions, such as uniform backgrounds and consistent lighting (e.g., scanned documents).
-However, these techniques may fall short when dealing with text-in-wild scenarios [128], where the background and lighting are highly diverse and the text orientation is arbitrary.
-For tasks involving the diverse image properties of natural scenes, such as identifying signs in urban areas [129], two primary subtasks are distinguished.
-Scene-text detection focuses on identifying and localizing text regions within an image without interpreting their content, while scene-text recognition converts these localized regions into machine-readable characters [130, 131, 132].
-The combination of these tasks is known as scene-text spotting, which handles text detection and recognition in an end-to-end manner.
-These methods are generally faster and more efficient due to their unified feature extraction module [133, 134].
-Focusing solely on character detection and recognition can be insufficient when image conditions are highly diverse, such as varying lighting and backgrounds.
-Relying on character detection in diverse scenes can increase false positives, as many patterns may look similar to letters.
-Additionally, DL methods require vast amounts of data, and character-level annotation is more expensive than word-level or line-level annotation due to its granularity.
-Table 2: Main approaches to ILU identification. The table reports how often each combination appears in the surveyed literature. $\bullet$ = detection, $\bullet$ = recognition.
-Object Character Text Count Percentage
-20 54.1%
-4 10.8%
-3 8.1%
-2 5.4%
-2 5.4%
-2 5.4%
-1 2.7%
-1 2.7%
-1 2.7%
-1 2.7%
-This shift from character detection and recognition to text detection and recognition is shown in Table 2, where the majority of articles employ DL methods for both scene-text detection and recognition tasks.
+Table 2 shows that the majority of articles employ DL methods for both scene-text detection and recognition tasks.
 The share of hybrid approaches using non-DL methods for one of these steps are minimal.
@@
 Instead, researchers largely favour more comprehensive pipelines that rely on DL in each step.
-This shift towards unified DL-based approaches for both scene-text detection and recognition highlights their potential.
-However, to further enhance efficiency, researchers should focus more on text spotting [136, 134].
-In ILU identification, text recognition remains the greatest challenge.
-Environmental conditions and technical issues such as unreliable hardware or connectivity problems between devices can cause delays at the gate.
-These delays increase waiting times for stakeholders, lead to congestion, add to environmental pollution, and create further setbacks for subsequent loading units.
-When recognition errors occur, the risk of incorrect inventory lists rises, which not only requires time-consuming manual corrections in the handling process but can also raise security concerns [14].
-To address these challenges more effectively, end-to-end approaches have been proposed.
-By combining detection and recognition into a single forward pass with a unified feature extraction module, they streamline the process and improve overall efficiency.
-Few studies introduce object detection as a prior step to scene-text detection and recognition [73, 80].
-Detecting the ILU first reduces the search area for the text detection step and enables association of detected and recognised IDs with other objects, thereby improving scene understanding [73].
+By 2021–2025, DL-based approaches had established themselves as the predominant paradigm for ILU identification.
+Even hybrid DIP-based methods virtually disappeared, indicating that the DL models now offer both the required accuracy and operational robustness [73, 129, 130].
+The decline of hybrid approaches suggests that most effective solutions now rarely require DIP or traditional ML methods, except in specialised scenarios such as real-time constraints [106, 13, 19, 127, 126, 12].
+Researchers also address real-time constraints by focusing on text spotting as these methods are generally faster and more efficient due to their unified feature extraction module [131, 132, 133].
+A few studies introduce object detection or instance segmentation as a prior step to text detection and recognition [17, 84].
+Detecting the ILU reduces the search area for the text detection step and enables association of detected and recognised IDs with other objects, thereby improving scene understanding [17].
 4.2 Datasets
-Datasets form the backbone of reliable ILU detection.
-The quality, diversity, and annotation of the underlying data strongly influence how well models perform once deployed in real-world port environments.
+The quality and diversity of annotated data strongly influence how well DL models perform once deployed in real-world port environments.
 Carefully curated datasets with consistent labelling not only enable robust training and evaluation, but also determine whether methods can generalise beyond laboratory settings.
 Understanding the properties of datasets is therefore essential for evaluating the performance and applicability of CV-based ILU identification systems.
-Below, we summarise the main characteristics and settings of the datasets used in the eligible studies, highlighting camera positioning, diversity, and availability.
-4.3.1 Image Acquisition Setups
+Below, we summarise the main characteristics of the datasets used in the eligible studies, highlighting camera position settings, diversity, and availability.
+4.2.1 Image Acquisition Settings
 We categorised cameras as fixed or mobile cameras based on their mounting and usage.
@@
 Figure 5: Sample images for fixed perspective (left) [79], ground vehicle perspective (middle) [17], aerial perspective (right) [15].
-In contrast, mobile cameras are mounted on unmanned aerial vehicles, reach stackers, terminal trucks, or other industrial vehicles.
-For ground-level perspectives, we also considered handheld images of ILUs, as they closely resemble those captured by vehicle-mounted systems.
+In contrast, mobile cameras are mounted on unmanned aerial vehicles, reach stackers, terminal trucks, or other industrial vehicles and smart devices.
+For ground-level perspectives, we also considered handheld cameras, as they closely resemble those captured by vehicle-mounted systems.
 Unlike fixed setups, mobile cameras provide flexibility by moving across the terminal and capturing a wider range of perspectives.
-This mobility enables real-time port monitoring and localisation of each ILU in real-world coordinates, as well as monitoring their positions by handling vehicles such as reach stackers [138, 139].
-While they cover wider spatial areas and provide richer scene diversity, mobile cameras introduce more complex image variability, requiring advanced processing methods and diverse publicly available datasets.
+This mobility introduce higher image variability, requiring advanced processing methods and diverse datasets.
+When these challenges addressed, there is a potential for more precise and real-time terminal monitoring.
+Thanks to sensor equipped vehicles [135, 83], CV pipelines beyond ILU identification can be created.
+For example, localisation of ILUs through pose estimation, autonomous path optimisation for vehicles navigating around terminals, multi-agent coordination (collaboration of aerial and ground vehicles) for terminal monitoring.
 Table 3: Summary of dataset characteristics: camera position and dataset availability.
 Category Count Percentage
-Camera position Only Fixed 47 74.60 %
-Fixed and Ground 10 15.87 %
-Fixed, Aerial and Ground 3 4.76 %
-Only Ground 1 1.58 %
-Only Aerial 1 1.58 %
-Not reported 1 1.58 %
-Availability Not public 54 85.71 %
-Fully available 3 4.76 %
-Partially available 3 4.76 %
-Upon request 3 4.76 %
-Behind paywall 1 1.58 %
-As shown in Table 3 the most common collection setting is the fixed CCTV cameras at OCR-gates [140, 123, 141] or cranes (74.60 %) [80, 108] and a mixture of these images with mobile phone snapshots (15.87 %) [67].
+Camera position Only Fixed 47 74,60 %
+Fixed and Ground 10 15,87 %
+Fixed, Aerial and Ground 3 4,76 %
+Only Ground 1 1,58 %
+Only Aerial 1 1,58 %
+Not reported 1 1,58 %
+Availability Not public 54 85,71 %
+Fully available 3 4,76 %
+Upon request 3 4,76 %
+Partially available 2 3,17 %
+Behind paywall 1 1,58 %
+As shown in Table 3 the most common collection setting is the fixed CCTV cameras at OCR-gates [136, 125, 137] or cranes (74,60 %) [84, 110] and a mixture of these images with smartphone images (15,87 %) [74].
 However, vehicle-mounted cameras are increasingly used alongside fixed cameras to capture wider areas and provide information across the terminal [14, 15, 16, 17].
 4.2.2 Diversity
-Most datasets consist of images gathered from the real-world settings without controlled environments such as the World’s busiest ports like Shanghai [16], Singapore [44], Hong Kong [143], Kaohsiung [13], and Hamburg [73], as well as smaller inland ports like Dresden and Riesa [84].
+Most datasets consist of images gathered from ports, reflecting their operational settings.
+Images in these datasets were taken from the world’s busiest ports like Shanghai [13], Singapore [42], Hong Kong [138], Kaohsiung [9], and Hamburg [17], as well as other ports in Malaysia [73], Sweden [125], Vietnam [19, 126, 127, 75], Thailand [21], South Korea [14, 78, 139, 110, 104], Japan [101], and Italy [140].
+While rare, there are datasets with images collected in smaller inland ports like Dresden and Riesa [15, 17].
 Besides the terminal-specific data collection, web scraping were also used [129, 79].
 In addition, controlled environments such as the Shanghai Maritime University test facility [40] were used for data collection.
-Other datasets were created with images taken in Malaysia [66], Sweden [123], Vietnam [18, 124, 127, 68], Thailand [20], South Korea [142, 71, 144, 108, 102], Japan [99], and Italy [145].
 However, datasets consisting of images from multiple locations are rare.
-For DL methods, capturing the nuances of ILUs across different countries and their respective ports can be important.
-The number of images used in the experiments of the eligible articles ranges from a minimum of 30 images to a maximum of 34000, with a median of 1050 images.
-Since the DIP-based methods does not require a huge amount of data, researchers who tackled this problem before DL, experimented with a lower number of images compared to recent studies.
+For DL methods, capturing the nuances of ILUs across different countries and their respective ports is important.
+The number of images used in the experiments in the eligible articles ranges from a minimum of 30 images to a maximum of 34.000, with a median of 1.050 images.
+Since the DIP-based methods do not require a large amount of data, researchers who tackled ILU identification task before DL, experimented with a lower number of images compared to recent studies.
 However, with the emergence of ML and DL, the need for data has increased significantly.
-Most of the studies have addressed this need by pre-training their models on general-purpose scene-text detection and recognition datasets such as ICDAR2015 [146], CTW1500 [147], and SynthText [148], then fine-tuning on smaller ILU datasets.
-Early datasets already included efforts to represent diverse conditions in order to evaluate generalisation.
+Most of the studies have addressed this need by pre-training their models on general-purpose scene-text detection and recognition datasets such as ICDAR2015 [141], CTW1500 [142], and SynthText [143], then fine-tuning on smaller ILU datasets [16, 17].
+Starting from the early datasets, efforts to represent diverse conditions have been made in order to evaluate generalisation.
 Lighting variability is the most discussed factor, with datasets often including images taken in bright, low-light, and uneven lighting scenarios [102, 41].
-Considering how the earlier, thresholding-based methods may get affected by small changes in light intensity, diversifying the lighting conditions in the experiments was a major focus [145, 97, 143, 96].
+Considering how the earlier, thresholding-based methods get affected by small changes in light intensity, diversifying the lighting conditions in the experiments was a major focus [140, 99, 138, 98].
 However, currently lighting conditions are classified with more general classes: Day and night settings are frequently included to test the adaptability of the systems [106, 16, 17, 13].
 Weather conditions such as rain, snow, and fog are also represented, as they can significantly affect image quality and ILU identification performance [17, 74, 130, 138].
-Additionally, some data sets feature images with motion blur, camera noise, dirty, rusty, damaged, or partially occluded ID codes to further challenge algorithms with the exceptional conditions that exist in real-world settings [104, 73, 105, 126, 72].
-Rarely were these conditions simulated to increase the data sets [125].
+Additionally, some datasets feature images with motion blur, camera noise, dirty, rusty, damaged, or partially occluded ID codes to further challenge algorithms with the conditions that exist in real-world settings [106, 17, 107, 130, 79].
+Rarely were these conditions simulated to increase the diversity of datasets [129].
 4.2.3 Availability
+We divided datasets into 5 different categories: Fully available [24, 107, 17], upon request [144, 16, 136], partially available [145, 79], behind paywall [14] and not public.
 Table 3 shows the availability of the datasets used by the identified articles.
 The significant majority of the datasets are not publicly available, which creates a substantial barrier for research and development in this area.
-The lack of availability also prevents the reproducibility of results and fair comparison of the proposed methods.
-Further details for the publicly available datasets are shown in Table 4.
+We were not able to confirm the availability of the datasets that claim to be available upon request as our requests went unanswered.
+Recent papers acknowledge the lack of benchmark datasets and its negative effects on the field [16, 17].
+The lack of availability prevents the reproducibility of results and fair comparison between the proposed methods for researchers.
+As for the practitioners, while large seaports typically have extensive propriety datasets, publicly available datasets extend these capabilities to smaller terminals, facilitating data-driven optimisation.
+After screening the literature, we could find only 3 publicly available datasets.
 Among these, the TRansportation Unit Detection and Identification (TRUDI) dataset [17] stands out for its annotation density.
 It offers 35,034 instances across 733 images, covering five classes: Container, trailer, tank container, ID code, and logo.
-The Shipping Container Code dataset [23] leads in the number of images, and the Container Number-OCR dataset [105], though smaller (2,851 images from Tianjin Port), offers similar challenging conditions as others: variable weather conditions and degraded ID codes (dirt, rust, damage).
-Uniquely, TRUDI introduces motion blur as an additional challenging condition, as most of its images were captured via UAVs, reach stackers, and terminal trucks.
+It introduces motion blur as an additional challenging condition, as most of its images were captured via UAVs, reach stackers, and terminal trucks.
+The Shipping Container Code dataset [24] leads in the number of images, and the Container Number-OCR dataset [107], though smaller (2.851 images from Tianjin Port), offers similar challenging conditions as others: variable weather conditions and degraded ID codes (dirt, rust, damage).
+Further details for the publicly available datasets are shown in Table 4.
+Adopting these publicly available datasets for benchmarking will be very helpful for both researchers and practitioners to evaluate and compare proposed methods.
 Table 4: Publicly available ILU identification datasets.
 Dataset Images Instances Conditions Location
-Shipping container code [23] 3600 4000 Day/night, Rainy, snowy Dirty, damaged ID Not reported
-Container Number-OCR [105] 2851 2851 Low/high light Foggy Dirty, damaged ID Tianjin Port
-TRUDI [73] 733 35034 Day/night Rainy, snowy, foggy Motion blur Dirty, damaged ID
+Shipping container code [24]
+3600
+4000
+Day/night, Rainy, snowy Dirty, damaged ID Not reported
+Container Number-OCR [107]
+2851
+2851
+Low/high light Foggy Dirty, damaged ID Tianjin Port
+TRUDI [17]
+733
+35034
+Day/night Rainy, snowy, foggy Motion blur Dirty, damaged ID
 Multiple countries,
@@
 As there are multiple stages of ILU identification using CV methods, such as text detection and text recognition, each stage can be evaluated individually.
-However, end-to-end accuracy or success rate, is the most relevant metric due to its practical importance.
+In the literature, 60.18 % articles reported the results of individual stages only and did not include end-to-end results.
+However, the results for individual stages offer limited insight for the applicability of the method for the ILU identification task.
+End-to-end accuracy or success rate is the most relevant metric due to its practical importance.
 It is the most strict metric as one misrecognised character is enough for failed identification.
-Since the ILU identification problem is mostly solved with multiple stages such as text detection and text recognition, some articles only report the results of individual stages and do not include end-to-end results.
-However, the results for individual stages offer limited insight for the applicability of the method for the ILU identification task.
 Precision and recall are the common metrics for these individual stages.
@@
 The F1-score, or H-mean, harmoniously combines precision and recall into a single metric, offering a balanced view of the model’s performance.
-For object detection or text detection tasks, mean average precision (mAP) and average precision (AP) are commonly used, measuring precision across various intersection over union (IoU) levels.
-The IoU metric is important for assessing the overlap between predicted and actual bounding boxes.
-As for the recognition, besides precision, recall and F1-score, other metrics such as the character error rate (CER) is used.
+For object detection, instance segmentation or text detection tasks, average precision (AP) and mean average precision (mAP) are commonly used, measuring precision across various intersection over union (IoU) levels.
+The IoU metric is important for assessing the overlap between predicted and ground truth bounding boxes.
+As for the recognition, besides precision, recall and F1-score, other metrics such as the character error rate (CER) or word accuracy are used.
 CER measures the proportion of characters that were incorrectly predicted compared to the total number of characters in the reference text, using the Levenshtein-distance [146].
@@
 Model complexity is assessed through the number of parameters, giga-floating point operations per seconds, and model size, indicating the computational demands and storage requirements.
-Although the recent studies use the same, established metrics adapted from general purpose text detection and recognition task, putting the reported results for ILU identification into a frame is not possible with the limited amount of publicly available datasets.
-Due to different datasets used in the experiments, the reported end-to-end accuracy of the proposed methods ranges from 5 % [142] to 96,17 % [144].
-This level of variance can be observed in every metric, including the processing speed.
+Although the recent studies use the same, established metrics adapted from general purpose text detection and recognition task, putting the reported results for ILU identification into a frame is not possible with the lack of an widely adopted benchmark dataset and the lack of end-to-end results.
+Only around 40 % of the studies reported end-to-end accuracy and it ranges from 5 % [14] to 99,51 % [74], with the mean of 84.19 % and standard deviation of 24.02 percentage points.
+Considering the high variation and the low number of reported end-to-end results ($n=25$), we did not perform further statistical analysis as these will not yield in statistically significant results.
 4.4 Future Research Directions
-In the literature, future directions for the ILU identification task mainly focus on enhancing the datasets by including more images taken from diverse environments such as various distances, angles, and under different weather and lighting conditions.
+In the literature, future directions for the ILU identification task mainly focus on enhancing the datasets by including more images taken from diverse environments, introducing different challenges such as various distances, angles, weather and lighting conditions.
 This expansion aims to cover multiple ports and scenarios to improve model generalisation.
 However, increasing the diversity of proprietary datasets would not solve the fundamental problem with the datasets for ILU identification.
-Recent papers acknowledge the lack of benchmark datasets and its negative effects on the field [122, 73].
 Given that creating a new dataset for downstream CV tasks are tedious, time consuming, and costly [147, 148], alternatives such as synthetic data generation can be considered [149, 143], e.g., by using a game engine [150].
-Handling challenging conditions, such as low-light scenarios, occlusions, and damaged ID codes, is another direction pointed by the researchers, with techniques proposed to manage partial occlusions and degraded text [105].
-Current solutions are relying on ”off-the-shelf” super-resolution methods such as Super-Resolution-CNN [154].
-Literature on scene-text super-resolution has focused primarily on natural language, leaving an opportunity for research specifically on contextless text such as ILU codes.
-Alternatively, multi-view recognition is also identified as a future direction due to its potential against degraded ID codes [23, 12] as multiple images of a single ILU are captured from different angles in this approach.
-Improving image quality through preprocessing is another focus area.
+Handling challenging conditions, such as low-light scenarios, occlusions, and degraded or damaged ID codes, is another direction pointed by the researchers.
 Suggested methods include image enhancement techniques such as noise reduction [18], or the use of Generative Adversarial Networks [151, 152] to address issues such as shadows, rust, and dirt [24, 107].
+Super-resolution [153] is a task that can tackle the problems with degraded or damaged ID codes.
+However, the literature on scene-text super-resolution has focused primarily on natural language, leaving an opportunity for research specifically on contextless text such as ILU codes.
+Alternatively, multi-view ILU identification is another future direction due to its potential against degraded ID codes [24, 8] as multiple images of a single ILU are captured from different angles in this approach.
 Postprocessing steps are also recommended to verify recognition results by using the check digit and the owner code.
 [12, 154, 76, 79].
-For DL-based methods, refining network architectures to increase the receptive field so that the model can see more of the scene in a single pass.
-It is important for detecting containers appears to be small due to perspective [122].
-Another future work suggestion is incorporating attention mechanisms to boost recognition accuracy [158].
-The development of unified, end-to-end models is suggested to streamline processing pipelines [73, 140, 105, 157].
+For DL-based methods, refining network architectures to increase the receptive field so that the model can see more of the scene, especially smaller objects in a single pass [16].
+Another future work suggestion is incorporating attention mechanisms to boost recognition accuracy [155] or optimising the tokenisation process of vision transformer-based architectures as there are always the same amount of characters expected for ISO6346-compliant ID codes.
+For better applicability, the development of unified, end-to-end models is suggested to streamline processing pipelines [17, 136, 107, 154].
+While DL models achieve high accuracy, their computational demands often limit deployment on edge devices like drones.
+Future work should focus on model compression, quantisation, or neural architecture search to enable real-time ILU identification in resource constrained environments.
 The contextless nature of ISO6346 codes makes scene-text spotting task harder as there are no natural language cues [156].
@@
 These new methods can also be useful for similar applications including International Union of Railways (UIC)-wagon codes, or registration plates.
-Improvements for real-time processing are also highlighted, with a focus on optimising algorithms for mobile devices and enhancing systems to work in real-time, potentially through video stream analysis [123, 42].
 5 Conclusion
-This review provides the first synthesis of CV-based ILU identification research by analysing 63 empirical studies published over three decades.
-There is a clear upward trend for the number of publications with the rise of DL methods and Asia’s dominance as a global trade hub, with 79.71 % of studies originating from the region.
-Public funding supports over a third of research highlighting a reliance on public investment for transportation related infrastructure.
-Policymakers could foster progress by funding collaborative initiatives and public-private partnerships, which currently account for just 6.35 % of studies.
-Although there are articles published in top journals and conferences, most studies were published in unranked outlets.
-The main reason for this is the direct application of general purpose CV methods to the ILU identification task.
-This lack of methodological contribution can be increased by adapting existing methods with modifications that addresses challenges, such as contextless text spotting and real-time processing constraints.
-Methodologically, the field has shifted from classical image processing to DL, currently used in more than 90 % of recent studies, thanks to GPU advancements and frameworks like PyTorch.
-This shift comes with the need of huge datasets to train DL models.
-Data limitations are the biggest obstacle with 85.71 % of datasets being private and most studies relying on controlled image acquisition setups.
-The median dataset size is too small for robust DL training, forcing researchers to adapt general text detection and recognition models.
-Another critical point for improving the quality and credibility of publications is using or developing benchmark datasets.
-The lack of benchmark dataset makes performance comparisons difficult, leaving gaps in understanding which methods work best in which setting.
-The research on ILU identification mainly focuses on fixed camera settings such as OCR-gates.
-However, recent studies show a shift from fixed to mobile cameras, such as drones or vehicle mounted cameras.
-This shift can spark more innovation due to new challenges related to dynamic scenes from the technical perspective [84] as well as social perspective such as privacy issues [161, 162, 163].
-Camera mounted vehicles offer new subtasks for the ILU identification task such as pose estimation of the ILUs by predicting the real-world coordinates (position) and heading (orientation).
-This way, a precise monitoring of the terminal is possible.
-To conclude, we suggest emphasising standardisation for evaluation with publicly available benchmark datasets, contextless text spotting architectures, real-time processing on edge devices and more focus on vehicle mounted camera setups for a more comprehensive ILU identification and terminal monitoring.
-Acknowledgements
-The project is supported by the German Federal Ministry for Digital and Transport (BMDV) in the funding program Innovative Hafentechnologien II (IHATEC II) with the funding number 19H23002C.
-Disclosure statement
-The authors report there are no competing interests to declare.
+This review provides the first systematic synthesis of CV-based ILU identification research by analysing 63 empirical studies published over three decades, and propose a more inclusive terminology.
+There is a clear upward trend for the quality and number of publications in the recent years driven by the adaptation of CV-based systems, which reduce human error and increase efficiency.
+Over the years, the field has shifted from classical DIP to DL thanks to the advancements in computational resources and deep learning frameworks.
+This shift comes with the need of huge datasets to train DL models and dataset accessibility is the biggest obstacle with 85,71 % of datasets being private.
+This creates problems for both researchers and practitioners.
+While large seaports have the means of creating large scale datasets, small sized ports have limitations to create datasets with the same scale, which creates a competitive disadvantage.
+The median dataset size is too small for robust DL training, forcing researchers to adapt general purpose text detection and recognition models, causing text recognition models to overfit on words with semantic meaning.
+Limited dataset accessibility also affects the quality and credibility of publications as there are no adapted benchmark datasets.
+This makes performance comparisons difficult along with the lack of end-to-end analysis, leaving gaps in understanding which methods perform best under which setting.
+To the best of our knowledge, there are only three publicly available datasets which can be adopted for benchmarking.
+Considering that two of them were published within 2025, in the near future, the adaptation of these datasets might resolve this problem.
+While the rise of DL-based methods brought the need for publicly available datasets, but the shift from static to dynamic image acquisition settings brings other dimensions.
+The research on ILU identification relied on fixed camera settings such as OCR-gates.
+However, recent studies are increasingly adopting vehicle mounted camera settings.
+This shift introduces new challenges like higher variations in object scale caused by changing camera-to-object distances and unknown ILU locations.
+This is why, other CV and robotics related tasks such as pose estimation (predicting the real-world coordinates and orientation of ILUs) or mission planning for autonomous vehicles (e.g.
+calculating the most optimal route for identifying ILUs) are now emerging priorities.
+The seemingly simple problem of identifying ILUs is a fundamental step that has not been taken for autonomous terminal operations.
+Therefore, addressing these challenges holds transformative potential for transportation and logistics as precise monitoring can optimise the terminal operations, reduce human errors, pave the way for faster turnaround times, lower operational costs and increased resilience against disruptions like misplaces ILUs.